# Conditionally valid Probabilistic Conformal Prediction

### v1

@@ Appendix A Additional results and calculations
$\mathbb{P}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\geq 1-\alpha-\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x}).$
@@
$\mathbb{P}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu)}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\geq 1-\alpha-\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x})\\ +\mathbb{P}\left({\varphi<f_{\tau_{x,z}}^{-1}\left({\lambda_{x,Y_{n+1},z}}\right)\leq Q_{1-\alpha}(\mu)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ -\mathbb{P}\left({Q_{1-\alpha}(\mu)<f_{\tau_{x,z}}^{-1}\left({\lambda_{x,Y_{n+1},z}}\right)\leq\varphi\,|\,X_{n+1}=x,Z_{n+1}=z}\right).$
A.2 Proof of Corollary 3.3
@@
$\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\mid X_{n+1}=x,Z_{n+1}=z}\right)\geq 1-\alpha-\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x})-p_{n+1}^{(x,z)},$
$p_{n+1}^{(x,z)}=\mathbb{P}\left({Q_{1-\alpha}(\mu)<f_{\tau_{x,z}}^{-1}\left({\lambda_{x,Y_{n+1},z}}\right)\leq\varphi\,|\,X_{n+1}=x,Z_{n+1}=z}\right),$
where $\mu=\frac{1}{n+1}\sum_{k=1}^{n}\delta_{f_{\bar{\tau}_{X_{k}}}^{-1}(\bar{\lambda}_{k})}+\frac{1}{n+1}\delta_{\infty}$. While $\mathbb{E}[p_{n+1}^{(X_{n+1},Z_{n+1})}]\leq\alpha$, studying $p_{n+1}^{(x,z)}$ is challenging. However, we control this term in Corollary A.5. Remark, the quantile $Q_{1-\alpha}(\mu)$ is an order statistic with a known distribution that rapidly converges to the true quantile $q_{0}$, which is defined for any $\epsilon\in[0,1-\alpha)$ by
$q_{\epsilon}=\inf\{t\in\mathbb{R}\colon\mathbb{P}(f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})\leq t)\geq 1-\alpha-\epsilon\}.$ (11)
Moreover, we define the cumulative density functions $F\colon t\mapsto\mathbb{P}(f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})\leq t)$ and $\hat{F}\colon t\mapsto\mathbb{P}(f_{\tau_{X,Z}}^{-1}(\lambda_{X,\hat{Y},Z})\leq t)$, where $(X,Y,Z)\sim\textup{P}_{X}\otimes\textup{P}_{Y\mid X}\otimes\Pi_{Z\mid X}$ and $(X,\hat{Y},Z)\sim\textup{P}_{X}\otimes\Pi_{Y\mid X}\otimes\Pi_{Z\mid X}$.
Corollary A.5.
Assume 1-2 hold, and let $x\in\mathbb{R}^{d}$ be such that $\Pi_{Y\mid X=x}$ is a probability measure. For any $\epsilon\in[0,1-\alpha)$, if $p_{\epsilon}=\mathbb{P}(f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})<q_{\epsilon})\leq 1-\alpha$, then it follows that
$p_{n+1}^{(x,z)}\leq\mathbb{P}\left({F^{-1}(1-\alpha-\epsilon)<f_{\tau_{x,y}}^{-1}(\lambda_{x,Y_{n+1},z})\leq\hat{F}^{-1}(1-\alpha)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ +\exp\left({-np_{\epsilon}(1-p_{\epsilon})h\left({\frac{1-\alpha-p_{\epsilon}}{p_{\epsilon}(1-p_{\epsilon})}}\right)}\right),$
@@
$A=\left\{{Q_{1-\alpha}(\mu)<q_{\epsilon}}\right\},$
$B_{x,z}=\left\{{y\in\mathcal{Y}\colon f_{\tau_{x,z}}(q_{\epsilon})<\lambda_{x,y,z}\leq f_{\tau_{x,z}}(\varphi)}\right\}.$
@@
$\mathbb{P}\left({f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu)}\right)<\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ \leq\mathbb{P}\left({A\mid X_{n+1}=x,Z_{n+1}=z}\right)+\mathbb{P}\left({Y_{n+1}\in B_{x,z}\mid X_{n+1}=x,Z_{n+1}=z}\right).$
@@
$\left\{{Q_{1-\alpha}(\mu)<q_{\epsilon}}\right\}\Leftrightarrow\left\{{\frac{1}{n+1}\sum_{k=1}^{n}\mathds{1}_{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})<q_{\epsilon}}\geq 1-\alpha}\right\}.$
@@
$\mathbb{P}\left({A\mid X_{n+1}=x,Z_{n+1}=z}\right)\leq\mathbb{P}\left({\sum_{k=1}^{n}\mathds{1}_{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})<q_{\epsilon}}\geq(n+1)(1-\alpha)}\right).$
Recall that $p_{\epsilon}=\mathbb{P}(f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})<q_{\epsilon})$, and also that we assume $p_{\epsilon}\leq 1-\alpha$. Therefore, the Bennett’s inequality [5, Theorem 2] implies that
$\mathbb{P}\left({A\mid X_{n+1}=x,Z_{n+1}=z}\right)\leq\exp\left({-np_{\epsilon}(1-p_{\epsilon})\,h\left({\frac{(n+1)(1-\alpha)-np_{\epsilon}}{np_{\epsilon}(1-p_{\epsilon})}}\right)}\right),$ (12)
@@
$\mathbb{P}\left({A\mid X_{n+1}=x,Z_{n+1}=z}\right)\leq\exp\left({-np_{\epsilon}(1-p_{\epsilon})h(u_{\epsilon})}\right).$
Furthermore, the definition of the cumulative distribution function:
$\mathbb{P}\left({Y_{n+1}\in B_{x,z}\mid X_{n+1}=x,Z_{n+1}=z}\right)\\ =\mathbb{P}\left({f_{\tau_{x,z}}(q_{\epsilon})<\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\mid X_{n+1}=x,Z_{n+1}=z}\right).$
By definition of $q_{\epsilon}$ provided in (11), we have $q_{\epsilon}=F^{-1}(1-\alpha-\epsilon)$. Moreover, for any $t\in(-\infty,\varphi)$, we have
$\hat{F}(t)$ $=\mathbb{P}\left({f_{\tau_{X,Z}}^{-1}(\lambda_{X,\hat{Y},Z})\leq t}\right)$
$=\int\mathbb{P}\left({f_{\tau_{X,Z}}^{-1}(\lambda_{X,\hat{Y},Z})\leq t\,\Big{|}\,X=x,Z=z}\right)\Pi_{Z\mid X=x}(\mathrm{d}z)\,\textup{P}_{X}(\mathrm{d}x)$
$=\int\mathbb{P}\left({\hat{Y}\in\mathcal{R}\left({x,f_{\tau_{z,z}}(t)}\right)\,\Big{|}\,X=x,Z=z}\right)\Pi_{Z\mid X=x}(\mathrm{d}z)\,\textup{P}_{X}(\mathrm{d}x).$
Using 2, the bijective property of $\tau\mapsto f_{\tau}(\varphi)$ implies the existence of $\nu\in\mathbb{R}$, such that $f_{\nu}(\varphi)=f_{\tau_{z,z}}(t)$. Note that, $\nu<\tau_{x,z}$ otherwise it would lead to $f_{\nu}(\varphi)\geq f_{\tau_{x,z}}(\varphi)>f_{\tau_{x,z}}(t)$. The definition of $\tau_{x,z}$ shows that
$\mathbb{P}\left({\hat{Y}\in\mathcal{R}\left({x,f_{\nu}(\varphi)}\right)\,\Big{|}\,X=x,Z=z}\right)<1-\alpha.$
Therefore, we deduce that $\hat{F}^{-1}(1-\alpha)\geq\varphi$, and we can conclude that
$\mathbb{P}\left({Y_{n+1}\in B_{x,z}\mid X_{n+1}=x,Z_{n+1}=z}\right)\\ \leq\mathbb{P}\left({F^{-1}(1-\alpha-\epsilon)<f_{\tau_{x,y}}^{-1}(\lambda_{x,Y_{n+1},z})\leq\hat{F}^{-1}(1-\alpha)\mid X_{n+1}=x,Z_{n+1}=z}\right).$ (13)
@@
$\epsilon_{n}=\sqrt{\frac{8\alpha(1-\alpha)\log n}{n}}.$
@@
If the distribution of $f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})$ is continuous, then for all $\epsilon\in[0,1-\alpha)$, we have $p_{\epsilon}=\mathbb{P}(f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})<q_{\epsilon})=1-\alpha-\epsilon$. Moreover, if $\epsilon_{n}\leq\frac{\alpha(1-\alpha)}{8}$, then it follows
@@
First, recall that $q_{\epsilon}$ is defined in (11). If the distribution of $f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})$ is continuous, then we have
$1-\alpha-\epsilon\leq F(q_{\epsilon})=\sup_{\delta>0}F(q_{\epsilon}-\delta)\leq\mathbb{P}\left({f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})<q_{\epsilon}}\right)=p_{\epsilon}\leq 1-\alpha-\epsilon.$

### v2

@@ Appendix A Additional Results and Calculations
    $\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\geq 1-\alpha-\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x}).$
    Combining the previous result with (10) and (11) shows that
    $\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu_{n})}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\geq 1-\alpha-\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x})\\ -\mathbb{P}^{\mathcal{T}}\left({Q_{1-\alpha}(\mu_{n})<f_{\tau_{x,z}}^{-1}\left({\lambda_{x,Y_{n+1},z}}\right)\leq\varphi\,|\,X_{n+1}=x,Z_{n+1}=z}\right).$
    Finally, using (9) gives a lower bound on $\mathbb{P}^{\mathcal{T}}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\mid X_{n+1}=x,Z_{n+1}=z}\right)$.
  2. 2.
    By definition of the total variation distance, we have
    $\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ \leq\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,\hat{Y}_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)+\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x}).$
    Moreover, Lemma A.3 implies that
    $\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,\hat{Y}_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)=\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau_{x,z}}(\varphi))).$
    Therefore, we deduce that
    $\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ \leq\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau_{x,z}}(\varphi)))+\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x}).$
@@
    $\mathbb{P}^{\mathcal{T}}\left({\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu_{n})}\right)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ \leq\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau_{x,z}}(\varphi)))+\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x})+q_{n+1}^{(x,z)}.$
A.2 Bound on $p_{n+1}^{(x,z)}$ and $q_{n+1}^{(x,z)}$
@@
$1-\alpha-\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x})-p_{n+1}^{(x,z)}\leq\mathbb{P}^{\mathcal{T}}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})\mid X_{n+1}=x,Z_{n+1}=z}\right)\\ \leq\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau_{x,z}}(\varphi)))+\mathrm{d}_{\mathrm{TV}}(\textup{P}_{Y\mid X=x};\Pi_{Y\mid X=x})+q_{n+1}^{(x,z)},$
In the following, we consider the cumulative density functions $F\colon t\mapsto\mathbb{P}^{\mathcal{T}}(f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})\leq t)$ and $\hat{F}\colon t\mapsto\mathbb{P}^{\mathcal{T}}(f_{\tau_{X,Z}}^{-1}(\lambda_{X,\hat{Y},Z})\leq t)$, where $(X,Y,Z)\sim\textup{P}_{X}\times\textup{P}_{Y\mid X}\times\Pi_{Z\mid X}$ and $(X,\hat{Y},Z)\sim\textup{P}_{X}\times\Pi_{Y\mid X}\times\Pi_{Z\mid X}$. We denote by $\mu$ and $\hat{\mu}$ the law of the random variables $f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})$ and $f_{\tau_{X,Z}}^{-1}(\lambda_{X,\hat{Y},Z})$. Moreover, recall that $\mu_{n}=\frac{1}{n+1}\sum_{k=1}^{n}\delta_{f_{\bar{\tau}_{X_{k}}}^{-1}(\bar{\lambda}_{k})}+\frac{1}{n+1}\delta_{\infty}$. Note, the quantile $Q_{1-\alpha}(\mu_{n})$ is an order statistic with a known distribution that converges to the true quantile $Q_{1-\alpha}(\mu)$. The quantile is defined for any $t\in(0,1)$ by
$Q_{t}(\nu)=\inf\{u\in\mathbb{R}\colon\nu((-\infty,u])\geq t\},\qquad\text{where $\nu\in\{\mu,\mu_{n},\hat{\mu}$\}}.$ (12)
Theorem A.5.
Assume 1-2 hold, and let $x\in\mathbb{R}^{d}$ be such that $\Pi_{Y\mid X=x}$ is a probability measure. For any $\epsilon\in[0,1-\alpha)$, if $p_{\epsilon}=\mathbb{P}^{\mathcal{T}}(f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})<Q_{1-\alpha-\epsilon}(\mu))\leq 1-\alpha$, then it follows that
$p_{n+1}^{(x,z)}\leq\mathbb{P}^{\mathcal{T}}\left({Q_{1-\alpha-\epsilon}(\mu)<f_{\tau_{x,y}}^{-1}(\lambda_{x,Y_{n+1},z})\leq Q_{1-\alpha}(\hat{\mu})\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ +\exp\left({-np_{\epsilon}(1-p_{\epsilon})h\left({\frac{1-\alpha-p_{\epsilon}}{p_{\epsilon}(1-p_{\epsilon})}}\right)}\right),$
@@
$A=\left\{{Q_{1-\alpha}(\mu_{n})<Q_{1-\alpha-\epsilon}(\mu)}\right\},$
$B_{x,z}=\left\{{y\in\mathcal{Y}\colon f_{\tau_{x,z}}(Q_{1-\alpha-\epsilon}(\mu))<\lambda_{x,y,z}\leq f_{\tau_{x,z}}(\varphi)}\right\}.$
@@
$\mathbb{P}^{\mathcal{T}}\left({f_{\tau_{x,z}}\left({Q_{1-\alpha}(\mu_{n})}\right)<\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\,|\,X_{n+1}=x,Z_{n+1}=z}\right)\\ \leq\mathbb{P}^{\mathcal{T}}\left({A\mid X_{n+1}=x,Z_{n+1}=z}\right)+\mathbb{P}^{\mathcal{T}}\left({Y_{n+1}\in B_{x,z}\mid X_{n+1}=x,Z_{n+1}=z}\right).$
@@
$\left\{{Q_{1-\alpha}(\mu_{n})<Q_{1-\alpha-\epsilon}(\mu)}\right\}\Leftrightarrow\left\{{\frac{1}{n+1}\sum_{k=1}^{n}\mathds{1}_{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})<Q_{1-\alpha-\epsilon}(\mu)}\geq 1-\alpha}\right\}.$
@@
$\mathbb{P}^{\mathcal{T}}\left({A\mid X_{n+1}=x,Z_{n+1}=z}\right)\leq\mathbb{P}^{\mathcal{T}}\left({\sum_{k=1}^{n}\mathds{1}_{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})<Q_{1-\alpha-\epsilon}(\mu)}\geq(n+1)(1-\alpha)}\right).$
Recall that $p_{\epsilon}=\mathbb{P}^{\mathcal{T}}(f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})<Q_{1-\alpha-\epsilon}(\mu))$, and also that we assume $p_{\epsilon}\leq 1-\alpha$. Therefore, the Bennett’s inequality (Boucheron et al., 2003, Theorem 2) implies that
$\mathbb{P}^{\mathcal{T}}\left({A\mid X_{n+1}=x,Z_{n+1}=z}\right)\leq\exp\left({-np_{\epsilon}(1-p_{\epsilon})\,h\left({\frac{(n+1)(1-\alpha)-np_{\epsilon}}{np_{\epsilon}(1-p_{\epsilon})}}\right)}\right),$ (13)
@@
$\mathbb{P}^{\mathcal{T}}\left({A\mid X_{n+1}=x,Z_{n+1}=z}\right)\leq\exp\left({-np_{\epsilon}(1-p_{\epsilon})h(u_{\epsilon})}\right).$
Furthermore, the definition of $B_{x,z}$ gives
$\mathbb{P}^{\mathcal{T}}\left({Y_{n+1}\in B_{x,z}\mid X_{n+1}=x,Z_{n+1}=z}\right)\\ =\mathbb{P}^{\mathcal{T}}\left({f_{\tau_{x,z}}(Q_{1-\alpha-\epsilon}(\mu))<\lambda_{x,Y_{n+1},z}\leq f_{\tau_{x,z}}(\varphi)\mid X_{n+1}=x,Z_{n+1}=z}\right).$
Moreover, for any $t\in(-\infty,\varphi)$, we have
$\hat{F}(t)$ $=\mathbb{P}^{\mathcal{T}}\left({f_{\tau_{X,Z}}^{-1}(\lambda_{X,\hat{Y},Z})\leq t}\right)$
$=\int\mathbb{P}^{\mathcal{T}}\left({f_{\tau_{X,Z}}^{-1}(\lambda_{X,\hat{Y},Z})\leq t\,\Big{|}\,X=x,Z=z}\right)\bar{\Pi}_{Z\mid X=x}(\mathrm{d}z)\,\textup{P}_{X}(\mathrm{d}x)$
$=\int\mathbb{P}^{\mathcal{T}}\left({\hat{Y}\in\mathcal{R}\left({x,f_{\tau_{z,z}}(t)}\right)\,\Big{|}\,X=x,Z=z}\right)\bar{\Pi}_{Z\mid X=x}(\mathrm{d}z)\,\textup{P}_{X}(\mathrm{d}x).$
Using 2, the bijective property of $\tau\mapsto f_{\tau}(\varphi)$ implies the existence of $\nu\in\mathsf{T}$, such that $f_{\nu}(\varphi)=f_{\tau_{z,z}}(t)$. Note that, $\nu<\tau_{x,z}$ otherwise it would lead to $f_{\nu}(\varphi)\geq f_{\tau_{x,z}}(\varphi)>f_{\tau_{x,z}}(t)$. The definition of $\tau_{x,z}$ shows that
$\mathbb{P}^{\mathcal{T}}\left({\hat{Y}\in\mathcal{R}\left({x,f_{\nu}(\varphi)}\right)\,\Big{|}\,X=x,Z=z}\right)<1-\alpha.$
Therefore, we deduce that $Q_{1-\alpha}(\hat{\mu})\geq\varphi$, and we can conclude that
$\mathbb{P}^{\mathcal{T}}\left({Y_{n+1}\in B_{x,z}\mid X_{n+1}=x,Z_{n+1}=z}\right)\\ \leq\mathbb{P}^{\mathcal{T}}\left({Q_{1-\alpha-\epsilon}(\mu)<f_{\tau_{x,y}}^{-1}(\lambda_{x,Y_{n+1},z})\leq Q_{1-\alpha}(\hat{\mu})\mid X_{n+1}=x,Z_{n+1}=z}\right).$ (14)
@@
$\epsilon_{n}=\sqrt{\frac{8\alpha(1-\alpha)\log n}{n}}.$ (15)
@@
If the distribution of $f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})$ is continuous, then for all $\epsilon\in[0,1-\alpha)$, we have $p_{\epsilon}=\mathbb{P}^{\mathcal{T}}(f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})<Q_{1-\alpha-\epsilon}(\mu))=1-\alpha-\epsilon$. Moreover, if $\epsilon_{n}\leq\frac{\alpha(1-\alpha)}{8}$, then it follows
@@
First, recall that $Q_{1-\alpha-\epsilon}(\mu)$ is defined in (12). If the distribution of $f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})$ is continuous, then we have
$1-\alpha-\epsilon\leq F(Q_{1-\alpha-\epsilon}(\mu))=\sup_{\delta>0}F(Q_{1-\alpha-\epsilon}(\mu)-\delta)\\ \leq\mathbb{P}^{\mathcal{T}}\left({f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})<Q_{1-\alpha-\epsilon}(\mu)}\right)=p_{\epsilon}\leq 1-\alpha-\epsilon.$